Can you show the how you worked out the problem and the formula you used?
In determining the relationship of a professor’s exams, a department checked students’ performances on the midterm and final exams in a class. Based on the following data, what can you say about the relationship of the exam scores? (a = 0.05)
H0:
H1:
a=
Decision Rule:
Student Midterm A 63
B 63 C 63 D 66 E 68 F 64 G 64 H 69
Test:
Final 110
118
103
120
125
125
120
120
Conclusion:
2b) Based on your results Midterm Exam score was 71?
above, what Final Exam score would you predict for a student whose
Answer #1

Solution :

a)

Given that =0.5

The hypothesis being tested is:

H0: There is no relationship between the test scores.  ß1=   0

Ha: There is a relationship between the test scores  ß1╪  0

Alpha = 0.05

Decision Rule: Reject H0 if p-value < 0.05.

Test:

x y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
63 110 4.0000 58.1406 15.250
63 118 4.0000 0.1406 -0.750
63 103 4.0000 213.8906 29.250
66 120 1.0000 5.6406 2.375
68 125 9.0000 54.3906 22.125
64 125 1.0000 54.3906 -7.375
64 120 1.0000 5.6406 -2.375
69 120 16.0000 5.6406 9.500
ΣX ΣY
Σ(x-x̅)²
Σ(y-ȳ)²
Σ(x-x̅)(y-ȳ)
Total sum 520.00 941.00 40.00 397.88 68.00
Mean 65.00 117.63 SSxx SSyy SSxy

sample size ,   n = 8          
here,

x̅ = Σx / n = 65.000   ,     ȳ = Σy/n = 117.625   
SSxx =    Σ(x-x̅)² =   40.0000          
SSxy=   Σ(x-x̅)(y-ȳ) =  68.0          
                  
estimated slope , ß1 = SSxy/SSxx = 68.0 / 40.000   = 1.70000

estimated std error of slope =Se(ß1) = Se/√Sxx =    6.859   /√   40.00  =  1.0845
                  
t stat = estimated slope/std error =ß1 /Se(ß1) = 1.7000 / 1.0845   = 1.5675
                  
t-critical value=    2.4469   [excel function: =T.INV.2T(α,df) ]          
Degree of freedom ,df = n-2=   6              
p-value = 0.1680              

The regression output is:

The p-value is 0.1680.

Since the p-value (0.1680) is greater than the significance level (0.05), we cannot reject the null hypothesis.

Conclusion :

Therefore, we have insufficient evidence to conclude that there is a relationship between the test scores.

b)

Final test score = 127.825.


Let me know in the comment section if anything is not clear. I will reply ASAP!

If you like the answer, please give a thumbs-up. This will be quite encouraging for me.Thank-you!